# Mitigating Content Effects on Reasoning in Language Models through Fine-Grained Activation Steering

### v1

@@ Abstract
Large language models (LLMs) frequently demonstrate reasoning limitations, often conflating content plausibility (i.e., material inference) with logical validity (i.e., formal inference). This can result in biased inferences, where plausible arguments are incorrectly deemed logically valid or vice versa. Understanding and mitigating this limitation is critical, as it undermines the trustworthiness and generalizability of LLMs in applications that demand rigorous logical consistency, such as legal reasoning, scientific analysis, and safety-critical decision support. This paper investigates the problem of mitigating content biases on formal reasoning through activation steering, an inference-time intervention technique that directly modulates internal model activations. Specifically, we first curate a controlled syllogistic reasoning dataset covering 24 types of logical argument schemes designed to disentangle formal validity from content plausibility. After localising the layers responsible for formal and material inference through probing, we investigate contrastive activation steering methods for test-time interventions. An extensive empirical analysis on different LLMs (i.e. Llama, Gemma, and Qwen) reveals that contrastive steering consistently supports linear control over content biases. However, we observe that a static steering approach is insufficient for achieving improvements on all the tested models. We then leverage the possibility to control content effects by dynamically determining the value of the steering parameters via fine-grained conditional methods. We found that conditional steering is effective in reducing biases on unresponsive models, achieving up to 15% absolute improvement in formal reasoning accuracy with a newly introduced kNN-based conditional method (K-CAST). Finally, additional experiments reveal that steering for content effects is robust to prompt variations, incurs minimal side effects on multilingual language modeling capabilities, and can partially generalize to out-of-distribution reasoning tasks. Practically, this paper demonstrates that activation-level interventions can offer a scalable inference-time strategy for enhancing the robustness of LLMs, contributing towards more systematic and unbiased formal reasoning [^1] Correspondence email: ac4mv@sheffield.ac.uk. Code and data will be made available upon publication..
@@ 1 Introduction
Large language models (LLMs) possess advanced natural and common-sense reasoning capabilities but are prone to content effects, i.e., systematic biases where prior knowledge and believability of content influence logical inference (bertolazzi-etal-2024-systematic, ; 10.1093/pnasnexus/pgae233, ). For example, an LLM may incorrectly judge a logically invalid syllogism as valid if its content aligns with common-sense knowledge (e.g., “All students read; some readers are professors; therefore some students are professors”), mirroring human content biases 10.1093/pnasnexus/pgae233 . Such behavior violates the requirement of formal reasoning (where validity should depend only on logical form, not content). Recent studies have documented these content-based reasoning failures in LLMs bertolazzi-etal-2024-systematic , showing that models, like humans, find factually believable premises easier to “prove” and struggle with abstract or counter-intuitive ones 10.1093/pnasnexus/pgae233 . This undermines LLMs reliability on formal reasoning tasks, particularly syllogistic logic-oriented tasks highlighted by bertolazzi2024systematic .
@@
On the other hand, prompting strategies alone are insufficient to eliminate content effects. Chain-of-thought (CoT) prompting wei2022chain and related methods (e.g. zero-shot CoT kojima2022large ) can improve reasoning performance by eliciting step-by-step logic. However, biases often persist in these generated explanations, and models may still arrive at content-biased conclusions even when “thinking aloud” ranaldi2025improving . In fact, bertolazzi2024systematic finds that while CoT-based in-context learning boosts accuracy on logical deductions, it does not fully remove biases like content believability effects. Similarly, straightforward instructions to “ignore prior knowledge” tend to be ineffective because the model’s latent representations already entangle content with reasoning dasgupta2022content . Supervised fine-tuning on reasoning data can mitigate some biases bertolazzi2024systematic , but it requires extensive data curation and does not explicitly guarantee content neutrality in the reasoning process. Similarly, neuro-symbolic approaches have been proposed to improve robustness in formal reasoning with LLMs quan-etal-2024-verification ; quan2025peirce ; pan-etal-2023-logic ; lyu-etal-2023-faithful . However, they typically introduce the complexity of integrating LLMs with external symbolic solvers through autoformalization wu2022autoformalization .
In this paper, we aim to go beyond these baselines by directly intervening internally into the model’s reasoning to enforce content-invariance (see Figure 1). We focus on syllogistic reasoning, a task that enables us to systematically disentangle formal reasoning from content by leveraging well-known logical-deductive argument schemes in the literature. Overall our contribution and findings can be summarised as follows:
@@
A large-scale dataset to disentangle formal and material inference.
@@
Expanding on previous work bertolazzi-etal-2024-systematic ; kim2024mechanistic ; wysocka-etal-2025-syllobio , we generate a synthetic dataset leveraging known syllogistic arguments considering the intersection of plausible/implausible and formally valid/invalid arguments. The dataset includes over 16k arguments generated by instantiating 24 abstract syllogistic schemes with the support of external knowledge bases (i.e., Wordnet miller1995wordnet ).
Probing and localizing formal and material inference,
@@
We perform an observational study through probing ferreira2021does ; belinkov2022probing to localise information about the validity and plausibility of arguments within the models. The experiments reveal that the information is maximally localised in later layers, peaking at the third quarter of the layers in the residual stream across different LLMs.
@@
Leveraging the observational study, we investigate static and contrastive activation steering methods panickssery2023steering . In general, we found that contrastive steering is effective on most of the tested models. In particular, the experiments reveal that steering vectors can explicitly control models’ output along a linear direction depending on the steering parameters, influencing the accuracy on both valid and invalid arguments. However, we found that static steering cannot improve performance on all the tested models.
@@
We adapt the recently proposed conditional activation steering (CAST) method mallen2024cast for content effects, and propose a new fine-grained variation employing a k-NN classifier to dynamically determine the steering parameters (K-CAST). We found that such methods can reduce biases on models that are unresponsive to static steering while, at the same time, increasing overall accuracy by up to $\approx$ 15% absolute value.
@@
We investigate the impact of steering for content effects on multilingual language modeling capabilities raffel2020exploring and out-of-distribution reasoning tasks PrOntoQA ; chan2024rulebreakers . We found that steering is well-localized, incurring minimal side effects on language modeling capabilities. At the same time, we found that steering vectors computed on the synthetic data can generalize to some extent to different reasoning tasks, with some variations across models. These results highlight both the potential of steering to improve targeted reasoning capabilities as well as the persisting challenges in enabling full generalization.
@@ 2 Background
Recent research has demonstrated that LLMs exhibit content effects in formal reasoning tasks, mirroring cognitive biases that may align or differ from those observed in humans dasgupta2022content ; kim2024mechanistic ; mondorf2024comparing ; seals2024evaluating ; wysocka-etal-2025-syllobio ; eisape2024systematic . These effects arise when the semantic plausibility of a prompt influences the model’s reasoning process, often leading to correct conclusions for plausible statements and systematic errors for implausible but logically valid ones bertolazzi-etal-2024-systematic .
dasgupta2022content first showed that LLMs perform better on reasoning tasks when the content of the problem aligns with world knowledge. In their experiments, models were significantly more accurate on syllogistic tasks when the conclusions were semantically plausible, even when this plausibility conflicted with the actual logical validity of the argument. This indicates a bias toward material reasoning (reasoning grounded in semantic associations) rather than formal reasoning (reasoning based strictly on logic).
Further work by bertolazzi-etal-2024-systematic systematically evaluated LLMs on a broad suite of syllogisms and found that performance dropped sharply for arguments that contradicted commonsense knowledge. This reliance on content plausibility suggests that LLMs are susceptible to semantic interference, failing to uphold the norms of formal logic when they conflict with prior knowledge. seals2023llms and agarwal2024mind also emphasized this discrepancy, showing that even the most capable frontier models tend to conflate logical validity with content-driven plausibility. To the best of our knowledge, this is the first work investigating how content effects can be reduced through activation steering techniques subramani2022extracting ; hernandezinspecting ; zou2023representation ; turner2023steering ; li2023inference ; zhao2024steering .
@@ 3 Methodology
Figure 1: Overview of our methodology for mitigating content effects on formal reasoning via activation steering. We first curate a controlled syllogistic reasoning dataset designed to disentangle formal validity from content plausibility. Subsequently, after localising the layers mostly responsible for formal and material inference through probing, we investigate static and conditional contrastive steering methods for test-time interventions to debias models’ behaviour.
Our goal is to investigate and mitigate content effects in LLMs, i.e., systematic biases where semantic plausibility influences logical reasoning. To this end, we design a controlled syllogistic reasoning task, leveraging 24 abstract syllogistic schemes automatically instantiated through taxonomic knowledge from external knowledge bases (Sec. 3.1) and apply activation-level steering techniques to modulate model behavior toward formal validity assessment (Sec. 3.2). Further, we identify the limitations of current state-of-the-art steering methods, and propose a more fine-grained steering approach (Sec. 3.2.1). Figure 1 provides a high-level overview of the methodology.
3.1 Formal Reasoning Task on Syllogistic Arguments
Inspired by recent work lample2019deep ; bertolazzi-etal-2024-systematic ; wysocka-etal-2025-syllobio ; kim2024mechanistic , we evaluate formal reasoning in LLMs through syllogistic arguments. We formalise the task of syllogistic reasoning as a binary classification problem, where the objective is to determine the formal validity of a syllogism, whether its conclusion follows logically from its premises, irrespective of the plausibility of the content. Specifically, the model is expected to predict VALID or INVALID based solely on the logical form in the syllogism structure $\mathcal{S}$.
@@ 3.1 Formal Reasoning Task on Syllogistic Arguments
Plausible Valid All apples are edible fruits. All edible fruits are fruits. All apples are fruits.
Implausible Valid All apples are vegetations. All vegetations are institutions. Some apples are institutions.
Plausible Invalid All apples are edible fruits. All apples are fruits. All edible fruits are fruits.
Implausible Invalid All apples are institutions. All vegetations are institutions. All apples are vegetations.
@@
We construct a dataset of approximately 16,000 syllogistic arguments in English to systematically analyze content effects. Each argument instantiates one of the 24 formal syllogistic schemas (see Appendix), ranging from categorical to disjunctive forms, and is explicitly varied along dimensions of formal validity and semantic plausibility.
Our data generation process begins with the formalization of syllogistic structures in first-order logic (FOL), following prior work on logical reasoning datasets (bertolazzi-etal-2024-systematic, ; wysocka-etal-2025-syllobio, ). For example, an AA1 schema is defined as:
$\forall x\ (A(x)\rightarrow B(x)),\quad\forall x\ (B(x)\rightarrow C(x))\Rightarrow\forall x\ (A(x)\rightarrow C(x))$
Each logical schema is then converted into natural language templates such as:
All A are B, All B are C, All A are C.
To control semantic content, we instantiate these syllogistic templates with concrete noun phrases drawn from WordNet [^2] https://wordnet.princeton.edu/miller1995wordnet based on its taxonomic hierarchies, using hypernym-hyponym relations.
@@ 3.2 Activation Steering
Activation steering is a causal intervention technique for modulating the internal computation of LLMs by linearly modifying hidden activations, also known as activation engineering subramani2022extracting ; hernandezinspecting ; zou2023representation ; turner2023steering ; li2023inference ; zhao2024steering . In this work, we adopt both static and conditional activation steering methods.
Contrastive Activation Steering (CAA)
[…]
The resulting steering direction is $\Delta\phi=\mu_{+}-\mu_{-}$. At inference time, we steer the model by modifying its internal activations:
$\tilde{\phi}(x)=\phi(x)+\alpha\cdot\Delta\phi$ (1)
where $\alpha$ is a scaling hyperparameter. This method has proven effective at shifting model outputs in desired directions with minimal degradation to fluency or relevance.
We apply CAA to reduce content effects by steering activations toward representations associated with content-invariant outputs. To this end, $\mu_{+}$ is computed from activations leading to correct formal validity predictions, while $\mu_{-}$ is computed from incorrect predictions affected by content bias.
@@
is a steering method designed to enable selective modulation of model behaviors by conditionally applying activation steering based on the input context mallen2024cast . Unlike traditional activation steering methods that uniformly apply steering vectors across all inputs, CAST introduces a mechanism to determine, at inference time, whether to apply a steering vector based on the similarity of the current input’s activation to predefined condition vectors.
Formally, let $\phi(x)\in\mathbb{R}^{d}$ denote the activation vector at a specified layer and position for input $x$. Given a labeled dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $y_{i}\in\{+1,-1\}$ denotes the presence or absence of a target condition on $x_{i}$ (i.e., the argument is formally valid), CAST computes a condition vector $\psi_{c}$ based on the average aggregation (or PCA) of individual activations vectors $\phi(x_{i})$ such that $y_{i}=+1$. During inference, for a given input $x$, the similarity between $\phi(x)$ and $\text{proj}\phi(x)_{\psi_{c}}$ – i.e., The projection of $\phi(x)$ onto $\psi_{c}$ – is computed, typically using cosine similarity:
$\text{sim}(\phi(x),\text{proj}\phi(x)_{\psi_{c}})=\frac{\phi(x)\cdot\text{proj}\phi(x)_{\psi_{c}}}{\|\phi(x)\|\|\text{proj}\phi(x)_{\psi_{c}}\|}$ (2)
@@
In this work, we adapt CAST to dynamically determine the value of the scaling parameter $\alpha$, since our empirical analysis on static contrastive steering reveals that the sign of $\alpha$ enables explicit control over the accuracy on valid and invalid arguments. In particular, given two condition vectors $\psi_{c_{+}}$ and $\psi_{c_{-}}$, the first computed for valid arguments and the second computed for invalid arguments, we modify the value of $\alpha$ dynamically according to the following function:
$f(\alpha,\phi(x),\psi_{c_{+}},\psi_{c_{-}})=\begin{cases}-\alpha&\text{if }\text{sim}(\phi(x),\text{proj}\phi(x)_{\psi_{c_{+}}})>\text{sim}(\phi(x),\text{proj}\phi(x)_{\psi_{c_{-}}})\\ \alpha&\text{otherwise}\end{cases}$ (3)
@@
3.2.1 K-CAST: kNN-Based Conditional Activation Steering

### v2

@@ Abstract
Large language models (LLMs) exhibit reasoning biases, often conflating content plausibility with formal logical validity. This can lead to wrong inferences in critical domains, where plausible arguments are incorrectly deemed logically valid or vice versa. This paper investigates how content biases on reasoning can be mitigated through activation steering, an inference-time technique that modulates internal activations. Specifically, after localising the layers responsible for formal and plausible inference, we investigate activation steering on a controlled syllogistic reasoning task, designed to disentangle formal validity from content plausibility. An extensive empirical analysis reveals that contrastive steering methods consistently support linear control over content biases. However, a static approach is insufficient to debias all the tested models. We then investigate how to control content effects by dynamically determining the steering parameters through fine-grained conditional methods. By introducing a novel kNN-based conditional approach (K-CAST), we demonstrate that conditional steering can effectively reduce biases on unresponsive models, achieving up to 15% absolute improvement in formal reasoning accuracy. Finally, we found that steering for content effects is robust to prompt variations, incurs minimal side effects on multilingual language modeling capabilities, and can partially generalize to different reasoning tasks. In practice, we demonstrate that activation-level interventions offer a scalable inference-time strategy for enhancing the robustness of LLMs, contributing towards more systematic and unbiased reasoning capabilities.
Code and datahttps://github.com/neuro-symbolic-ai/steering˙content˙effects
@@ 1 Introduction
Large language models (LLMs) possess advanced natural and common-sense reasoning capabilities but are prone to content effects, i.e., systematic biases where prior knowledge and believability of content influence logical inference (Bertolazzi et al. 2024a; Lampinen et al. 2024). For example, an LLM may incorrectly judge a logically invalid syllogism as valid if its content aligns with common-sense knowledge (e.g., “All students read; some readers are professors; therefore some students are professors”), mirroring human content biases (Lampinen et al. 2024). Such behaviour violates the requirement of formal reasoning, where validity should depend only on logical form, not content. Recent studies have documented these content-based reasoning failures (Bertolazzi et al. 2024a), showing that models find factually believable premises easier to “prove” and struggle with abstract or counter-intuitive ones (Lampinen et al. 2024). This undermines LLMs’ reliability on formal reasoning, particularly logic-oriented tasks highlighted by Bertolazzi et al. (2024b).
@@
On the other hand, prompting strategies alone are insufficient to eliminate content effects. Chain-of-thought (CoT) prompting and related methods (Wei et al. 2022; Kojima et al. 2022) can improve reasoning. However, biases often persist in the generated explanations, and models may still arrive at content-biased conclusions even when “thinking aloud” (Ranaldi et al. 2025). In fact, (Bertolazzi et al. 2024b) finds that while CoT and fine-tuning boost accuracy on logical deductions, they do not fully remove biases like content believability effects. Similarly, neuro-symbolic approaches have been proposed to improve robustness in formal reasoning with LLMs (Quan et al. 2024, 2025b, 2025a; Pan et al. 2023; Lyu et al. 2023). However, they introduce the complexity of integrating LLMs with external symbolic solvers.
Figure 1: Overview of our methodology for mitigating content effects on reasoning via activation steering. We first curate a controlled syllogistic reasoning dataset designed to disentangle formal validity from content plausibility. Subsequently, after localising the layers mostly responsible for formal and plausible inference through probing, we investigate static and conditional steering methods to debias models’ behaviour.
Unlike existing methods, we directly manipulate internal activations to investigate if content-invariance can be effectively enforced through test-time interventions and to gain a deeper understanding of the internal representational mechanisms (see Figure 1). Overall, our contribution and findings can be summarised as follows:
@@
A large-scale dataset to disentangle content from formal reasoning.
@@
Expanding on previous work (Bertolazzi et al. 2024a; Kim et al. 2025; Wysocka et al. 2025), we generate a synthetic dataset leveraging known syllogistic arguments, considering the intersection of plausible/implausible and formally valid/invalid arguments. The dataset includes over 16k arguments generated by instantiating 24 abstract syllogistic schemes with the support of Wordnet (Miller 1995).
Localizing formal and plausible inference.
@@
We perform an observational study through probing (Ferreira et al. 2021; Belinkov 2022) to localise information about the validity and plausibility of arguments within the models. The experiments reveal that the information is maximally localised in later layers, peaking at the third quarter of the layers in the residual stream across different LLMs.
@@
Leveraging the observational study, we investigate static and contrastive activation steering methods (Panickssery et al. 2023). In general, we found that contrastive steering is effective on most of the tested models. In particular, the experiments reveal that steering vectors can explicitly control models’ output along a linear direction depending on the steering parameters, influencing the accuracy on both valid and invalid arguments. However, we found that static steering cannot improve performance on all the tested models.
@@
We adapt the recently proposed conditional activation steering (CAST) method (Lee et al. 2025) for content effects, and propose a new fine-grained variation employing a k-NN classifier to dynamically determine the steering parameters (K-CAST). We found that such methods can reduce biases on models that are unresponsive to static steering while, at the same time, increasing overall accuracy by up to $\approx$ 15% absolute value.
@@
We investigate the impact of steering for content effects on multilingual language modeling capabilities (Raffel et al. 2020) and out-of-distribution reasoning tasks (Saparov and He 2023; Chan et al. 2024). We found that steering is well-localized, incurring minimal side effects on language modeling capabilities. At the same time, we found that steering vectors computed on the synthetic data can generalize to some extent to different reasoning tasks, with some variations across models. These results highlight both the potential of steering to improve targeted reasoning capabilities as well as the persisting challenges in enabling full generalization.
@@ 2 Background
Recent research has demonstrated that LLMs exhibit content effects in formal reasoning tasks, mirroring cognitive biases that may align or differ from those observed in humans (Dasgupta et al. 2022; Kim et al. 2025; Mondorf and Plank 2024; Seals and Shalin 2024; Wysocka et al. 2025; Eisape et al. 2024). These effects arise when the semantic plausibility of a prompt influences the model’s reasoning process, often leading to correct conclusions for plausible statements and systematic errors for implausible but logically valid ones (Bertolazzi et al. 2024a).
(Dasgupta et al. 2022) first showed that LLMs perform better on reasoning tasks when the content of the problem aligns with world knowledge. In their experiments, models were significantly more accurate on syllogistic tasks when the conclusions were semantically plausible, even when this plausibility conflicted with the actual logical validity of the argument. This indicates a bias toward material reasoning (reasoning grounded in semantic associations) rather than formal reasoning (reasoning based strictly on logic).
Further work by (Bertolazzi et al. 2024a) systematically evaluated LLMs on a broad suite of syllogisms and found that performance dropped sharply for arguments that contradicted commonsense knowledge. This reliance on content plausibility suggests that LLMs are susceptible to semantic interference, failing to uphold the norms of formal logic when they conflict with prior knowledge. (Seals et al. 2023) and (Agarwal et al. 2024) also emphasized this discrepancy, showing that even the most capable frontier models tend to conflate logical validity with plausibility. To the best of our knowledge, this is the first work investigating how content effects can be reduced through activation steering techniques (Subramani et al. 2022; Hernandez et al. 2024; Zou et al. 2023; Turner et al. 2023; Li et al. 2023; Zhao et al. 2024).
@@ 3 Methodology
Our goal is to investigate and mitigate content effects in LLMs, i.e., systematic biases where semantic plausibility influences logical reasoning. To this end, we design a controlled syllogistic reasoning task, leveraging 24 abstract syllogistic schemes automatically instantiated through taxonomic knowledge from external knowledge bases (Sec. 3.1) and apply activation-level steering techniques to modulate model behavior toward formal validity assessment (Sec. 3.2). Further, we identify the limitations of current state-of-the-art steering methods, and propose a more fine-grained steering approach (Sec. 3.2). Figure 1 provides a high-level overview of the methodology.
3.1 Formal Syllogistic Reasoning
Inspired by recent work (Lample and Charton 2019; Bertolazzi et al. 2024a; Wysocka et al. 2025; Kim et al. 2025), we evaluate formal reasoning in LLMs through syllogistic arguments. We formalise the task of syllogistic reasoning as a binary classification problem, where the objective is to determine the formal validity of a syllogism, whether its conclusion follows logically from its premises, irrespective of the plausibility of the content. Specifically, the model is expected to predict VALID or INVALID based solely on the logical form in the syllogism structure $\mathcal{S}$.
@@ 3.1 Formal Syllogistic Reasoning
  • Plausible Valid: All apples are edible fruits. All edible fruits are fruits. All apples are fruits.
  • Implausible Valid: All apples are vegetations. All vegetations are institutions. Some apples are institutions.
  • Plausible Invalid: All apples are edible fruits. All apples are fruits. All edible fruits are fruits.
  • Implausible Invalid: All apples are institutions. All vegetations are institutions. All apples are vegetations.
@@
We construct a dataset of approximately 16,000 syllogistic arguments in English to systematically analyze content effects. Each argument instantiates one of the 24 formal syllogistic schemas (details in the supplementary material) and is explicitly varied along dimensions of formal validity and semantic plausibility.
The data generation process begins with the formalization of syllogistic structures in first-order logic (FOL), following prior work on logical reasoning datasets (Bertolazzi et al. 2024a; Wysocka et al. 2025). Each logical schema is then converted into natural language templates such as: All A are B, All B are C, All A are C.
To control semantic content, we instantiate these syllogistic templates with concrete noun phrases drawn from WordNet [^1] https://wordnet.princeton.edu/(Miller 1995) based on taxonomic hierarchies, using hypernym-hyponym relations between concepts.
Figure 2: Linear probing results for formal validity on Llama-3.1 8b. The probing experiments reveal that the information about validity is maximally localised in the residual stream of later layers across different LLMs.
Model Size Base Model ($\alpha=0$) Steered Model ($\alpha_{best}$) $\mathbf{\alpha_{\text{best}}}$ $\Delta_{abs}^{Acc/CE}$ $\Delta_{rel}^{Acc/CE}$
Acc $\uparrow$ CE $\downarrow$ Acc/CE $\uparrow$ Acc $\uparrow$ CE $\downarrow$ Acc/CE $\uparrow$
Zero-shot
Llama 3.2 1b 58.17 44.04 1.32 73.56 6.35 11.58 -0.9 $\mathbf{10.26}$ $\mathbf{777.27}$
3b 77.79 17.50 4.45 77.79 17.50 4.45 0.0 $0.00$ $0.00$
Llama 3.1 8b 78.27 30.77 2.54 85.10 14.04 6.06 0.9 $3.52$ $138.58$
Gemma 2 2b 73.27 32.43 2.26 74.13 20.83 3.56 1.8 $1.30$ $57.52$
9b 85.00 8.46 10.05 83.27 1.92 43.37 0.6 $\mathbf{33.32}$ $\mathbf{331.54}$
Qwen 2.5 1.5b 75.67 14.42 5.25 77.79 12.88 6.04 0.3 $0.79$ $15.05$
3b 85.29 7.12 11.99 85.29 7.12 11.99 0.0 $0.00$ $0.00$
7b 88.85 5.39 16.48 89.90 0.96 93.65 -1.5 $\mathbf{77.17}$ $\mathbf{468.26}$
Few-shot
Llama 3.2 1b 57.21 45.70 1.25 66.44 6.08 4.94 -1.5 $\mathbf{3.69}$ $\mathbf{295.20}$
3b 72.79 28.14 2.59 78.27 22.31 3.51 0.3 $0.92$ $35.52$
Llama 3.1 8b 40.58 20.26 2.00 30.67 14.81 2.07 0.3 $0.07$ $3.50$
Gemma 2 2b 69.42 14.23 4.88 70.00 12.31 5.69 -0.3 $0.81$ $16.60$
9b 84.61 15.25 5.54 80.38 3.33 24.14 0.3 $\mathbf{18.60}$ $\mathbf{335.74}$
Qwen 2.5 1.5b 51.92 65.12 0.80 72.69 32.18 2.26 -1.2 $1.46$ $182.50$
3b 86.44 13.91 6.21 86.63 4.80 18.02 0.6 $\mathbf{11.81}$ $\mathbf{190.18}$
7b 89.80 9.36 9.60 89.90 4.81 18.70 0.9 $9.10$ $94.79$
Table 1: Results of contrastive steering with static values of $\alpha$. The table compares the performance of unsteered base models ($\alpha=0$) against the optimal steered performance ($\alpha_{\text{best}}$) with values of $\alpha$ selected from the interval [3.0,3.0][-3.0,3.0]. Acc/CE is the composite metric (Accuracy/Content Effect). The final columns quantify the Acc/CE gain: $\Delta_{\text{abs}}$ is the raw increase, and $\Delta_{\text{rel}}$ is the percentage increase relative to the baseline. The results reveal that contrastive steering is highly effective for most models, except Llama 3.2 3b and Qwen 2.5 3b.
@@ 3.2 Activation Steering
Activation steering is a causal intervention technique for modulating the internal computation of LLMs by linearly modifying hidden activations, also known as activation engineering (Subramani et al. 2022; Hernandez et al. 2024; Zou et al. 2023; Turner et al. 2023; Li et al. 2023; Zhao et al. 2024). In this work, we adopt both static and conditional activation steering methods.
Contrastive Activation Addition (CAA)
computes the steering vector using a set of labelled examples, based on the observed model behavior  (Panickssery et al. 2023).
Let $\phi(x)\in\mathbb{R}^{d}$ denote the activation vector at a chosen layer and token position for input $x$. Given a set of $N$ contrastive pairs of activations $\mathcal{P}=\{(a_{i}^{+},a_{i}^{-})\}_{i=1}^{N}$, where $a_{i}^{+}=\phi(x_{i}^{+})$ and $a_{i}^{-}=\phi(x_{i}^{-})$ are positive and negative activation vectors derived from inputs $x_{i}^{+}$ and $x_{i}^{-}$ leading to desired and contrasting behaviors respectively. The resulting steering vector is the mean difference between positive and negative activations:
$\Delta\phi=\frac{1}{N}\sum_{i=1}^{N}\left(a_{i}^{+}-a_{i}^{-}\right)$ (1)
At inference time, given a new input $x$, the model is steered by modifying its internal activations $\tilde{\phi}(x)=\phi(x)+\alpha\cdot\Delta\phi$, where $\alpha$ is a scaling hyperparameter determining the strength of the intervention.
We apply CAA to reduce content effects by steering activations toward representations associated with content-invariant outputs. To this end, the positive vectors are collected from inputs leading to correct formal validity predictions, while the negative vectors are collected from incorrect predictions affected by content bias.
@@
is a steering method designed to enable selective modulation of model behaviors by conditionally applying activation steering based on the input context (Lee et al. 2025). Unlike traditional activation steering methods that uniformly apply steering vectors across all inputs, CAST introduces a mechanism to determine, at inference time, whether to apply a steering vector based on the similarity of the current input’s activation to predefined condition vectors.
Formally, let $\phi(x)\in\mathbb{R}^{d}$ denote the activation vector at a specified layer and position for input $x$. Given a labeled dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $y_{i}\in\{+1,-1\}$ denotes the presence or absence of a target condition on $x_{i}$ (i.e., the argument is formally valid), CAST computes a condition vector $\psi_{c}$ based on the average aggregation (or PCA) of individual activations vectors $\phi(x_{i})$ such that $y_{i}=+1$. During inference, for a given input $x$, the similarity between $\phi(x)$ and $\pi_{\psi_{c}}(\phi(x))$ – i.e., the projection of $\phi(x)$ onto $\psi_{c}$ – is computed:
$\text{sim}(\phi(x),\pi_{\psi_{c}}(\phi(x)))=\frac{\phi(x)\cdot\pi_{\psi_{c}}(\phi(x))}{\|\phi(x)\|\|\pi_{\psi_{c}}(\phi(x))\|}$ (2)
@@
In this work, we adapt CAST to dynamically determine the value of the scaling parameter $\alpha$, since our empirical analysis on static steering reveals that the sign of $\alpha$ enables explicit control over the accuracy on valid and invalid arguments. In particular, given two condition vectors $\psi_{c_{+}}$ and $\psi_{c_{-}}$, the first computed for valid arguments and the second computed for invalid arguments, we modify the value of $\alpha$ dynamically according to the following function:
$\begin{split}f(\alpha,\phi(x),\psi_{c_{+}},\psi_{c_{-}})=\\ \begin{cases}-\alpha&\text{if }\text{sim}(\phi(x),\pi(\phi(x)_{\psi_{c_{+}}}))>\text{sim}(\phi(x),\pi(\phi(x)_{\psi_{c_{-}}}))\\ \alpha&\text{otherwise}\end{cases}\end{split}$ (3)
@@
K-CAST: kNN-Based Conditional Activation Steering
@@
(a) Llama 3b (Static)
(b) Llama 3b (K-CAST)
Figure 3: (a) Results of steering with static values of $\alpha$ on Llama 3b (note that $\alpha=0$ represents the performance of the base model without steering). (b) Impact of conditional steering (K-CAST) on Llama 3b. K-CAST increases accuracy while reducing content effects on models that are unresponsive to a static steering approach.